A butcher has some hamburger that is 22% fat and some hamburger that is 11% fat. How many pounds of each should be mixed to make 110 lb of hamburger that is 19% fat
Answer by richwmiller(17219)

a+b=110,
0.22*a+0.11*b=0.19*110
a=80, b=30